# Diversity Boosts AI-Generated Text Detection

### v2

@@ 1 Introduction
Contributions. We introduce DivEye  [^1] The code of our method and experiments is available at https://github.com/IBM/diveye/., a lightweight classifier trained on features extracted from off-the-shelf LLMs in a zero-shot manner. These features capture diversity-based statistics of token-level surprisal, which we leverage to improve AI-text detection. Our approach focuses on capturing the distributional irregularities in AI-generated text that arise from differences in the generative process compared to human writing.
@@ 3.1 Design Hypothesis
Our hypothesis is that human-written text inherently exhibits greater stylistic diversity and unpredictability than AI-generated text. In everyday writing, humans make creative, spontaneous choices, sometimes using unexpected words or phrases, that introduce bursts of surprise amid more routine language. Our approach centers on the premise that AI-generated text, despite its fluency, often lacks the inherent diversity observed in human-written language. This divergence stems from the fundamental objective of LLMs: to maximize the likelihood of generated sequences within their learned probability distributions (Park & Choi, 2024). Consequently, AI-generated text tends to exhibit a higher degree of predictability, resulting in lower variability and surprisal compared to human-authored content. We support this hypothesis through both intuitive reasoning and empirical evidence, as detailed in Remark A.
@@
Figure 2: Distributions of predicted class probabilities for diverse AI-text detectors. Trained and evaluated on Testbed 4 of the MAGE benchmark, DivEye shows stronger separation between Label 0 (human-written) and Label 1 (AI-generated), indicating greater confidence and discriminative power.
@@ 3.2 Mathematical underpinning of DivEye
Given a text sequence $X=\{x_{1},x_{2},\dots,x_{n}\}$, surprisal measures how ”unexpected” each token is in context. It can be computed directly from a model’s log-probabilities, providing a principled way to quantify the local unpredictability of text.
@@
$\mathcal{D}=\{\underbrace{\mu_{s},\sigma_{s}^{2},\gamma_{1},\gamma_{2}}_{\text{Distribution}}\oplus\underbrace{\Delta\mu,\Delta\sigma^{2}}_{\text{1st-Order}}\oplus\underbrace{\sigma^{2}_{\Delta^{2}},H_{\Delta^{2}},\rho_{\Delta^{2}}}_{\text{2nd-Order}}\}$ (6)
@@
DivEye as a booster. Existing detectors, whether fine-tuned classifiers or zero-shot LLM-based methods, primarily rely on semantic or surface-level cues, and often falter against high-quality adversarial examples that closely mimic human writing. DivEye offers a complementary signal by capturing statistical and temporal patterns of token-level unpredictability that are orthogonal to traditional features. We integrate DivEye into both settings by augmenting detector outputs with its feature vector and training a lightweight meta-classifier (e.g., XGBoost (Chen & Guestrin, 2016) or Random Forest (Breiman, 2001)) over the combined representation. Empirically, we find that this fusion significantly boosts performance, particularly on adversarial and out-of-distribution examples, without requiring retraining or modification of the original model.
@@ 4.1 Experimental Settings
Implementation Details & Metrics. Unless stated otherwise, we use GPT-2 to compute all DivEye feature vectors. Regardless, we studied the effect of DivEye with different LLMs as base models and summarized the results in Section 4.6. In score-only detection scenarios, predictions are based solely over concatenated DivEye features. For both standalone and boosted setups, we train a lightweight XGBoost (Chen & Guestrin, 2016) classifier as a meta-model, using only DivEye features in the former, and concatenating them with the original detector’s prediction scores in the latter. Each testbed in MAGE & RAID provides predefined training and test sets, which we use for model training and evaluation. We evaluate all models using Average Accuracy (AvgAcc), AUROC, and F1 score to capture overall, threshold-independent, and balanced performance, respectively.
@@ 4.2 Performance of DivEye
We evaluate DivEye across a wide range of challenging testbeds to assess its robustness and adaptability to both domain and model shifts. Table 1 presents the performance of DivEye on six distinct testbeds from the MAGE benchmark (Li et al., 2024): three in-distribution and three out-of-distribution. Across all testbeds, DivEye consistently achieves superior AUROC of $0.92$ on average and AvgAcc compared to existing zero-shot baselines, showcasing its ability to generalize effectively to both seen and unseen generation settings. We demonstrate that human-written and machine-generated text can be distinguished based on the hypothesis outlined in Section 3.
Table 4.2 benchmarks DivEye on the RAID dataset (Dugan et al., 2024), which includes a suite of diverse models, domains, attacks, and decoding strategies. DivEye outperforms a diverse set of zero-shot methods by $13.73\%$ and matches the performance of generative detection baselines, reaffirming its robustness to evasive generation strategies. Figures 3, 6 & 7 demonstrate the performance of DivEye across different domains and generator models, achieving competitive AUROC of $0.98$ and $0.93$, respectively. These results highlight DivEye’s stability and high performance across heterogeneous scenarios, underscoring its domain and model-agnostic nature.
@@
Table 2: Performance of zero-shot and open-source finetuned methods on RAID. Results are aggregated over 8 domains, 12 models, and 4 decoding strategies. $\delta$ denotes difference in AvgAcc from benchmark leader.
Frameworks Type AvgAcc $\delta$
Desklib AI [27] Fine-tuned 94.9% 0%
e5-small-lora [29] Fine-tuned 93.9% -1%
DivEye (Ours)
Classifier
(zero-shot)
93.63% -1.27%
Binoculars [45] Zero-shot 79.0% -15%
SuperAnnotate [89] Fine-tuned 70.3% -24.6%
RADAR [47] Fine-tuned 65.6% -29.3%
GLTR [38] Zero-shot 59.7% -35.2%
@@ 4.3 Robustness to Adversarial Attacks and Multilingual Text
Table 3: Performance of DivEye and baselines on adversarial benchmarks, MAGE & RAID. The RAID benchmark, which independently tests each model, does not report an AUROC score.
Settings Methods AvgAcc AUROC [MAGE] Testbed 8: Paraphrasing Attack Paraphrased via GPT-3.5-Turbo Longformer [10] 69.34% 0.76 BiScope [44] 69.30% 0.81 DivEye (Ours) 76.49% 0.87 [RAID] Adversarial Attacks Paraphrase, Whitespace, Misspelling, Homoglyph, Article Deletion & more Desklib AI [27] 91.2% - e5-small-lora [29] 85.7% - DivEye (Ours) 80.52% - Binoculars [45] 69.32% - RADAR [47] 63.9% - GLTR [38] 51.5% -
@@ 4.5 Effectiveness of Boosting by DivEye
We empirically verify that DivEye-based diversity features can act as performance boosters for a wide range of detection models. To integrate DivEye, we concatenate its feature vector with the original model’s prediction scores and train a lightweight XGBoost classifier as a meta-model. Table 4 illustrates improvements in AUROC and AvgAcc when diversity metrics are appended to existing frameworks such as RADAR, Binoculars, DetectLLM, BiScope and FastDetectGPT. Across all evaluated baselines, the inclusion of diversity features consistently leads to better detection scores by over $18.7\%$. Additionally, existing frameworks in combination with DivEye demonstrate substantial performance gains when evaluated against paraphrasing attacks. This validates the hypothesis that static and dynamic surprisal-based features capture orthogonal information to traditional heuristics, making them a valuable addition to any detection pipeline. We further explore the relative importance of DivEye and the base detector during prediction in Appendix D.4.
@@ 4.6 Ablation Studies
DivEye’s Performance on Different Base Models. To assess the adaptability of DivEye across different language model backbones, we evaluate its performance, on Testbed 4 of the MAGE benchmark, when instantiated with various base LLMs used to compute token-level surprisal. As shown in Figure 5(a), DivEye consistently performs well across all models, achieving an AUROC of $0.88$ with GPT2, $0.91$ with Llama-3.1-8B, and $0.90$ with Falcon-7B. Notably, even the smallest model, GPT2, performs competitively, and human classification accuracy improves with larger models, suggesting that higher-capacity LMs better capture stylistic diversity. These results highlight DivEye’s robustness and efficiency across scales, making it suitable for resource-constrained settings. Appendix D.3 further reports baseline performance across various base models.
Relevance of DivEye’s Features. DivEye’s feature set (Equation 6) captures token-level surprisal patterns across multiple orders, including distributional moments, first-order shifts, and second-order dynamics. To assess the contribution of each group, we compute feature importances from a trained XGBoost model. On average, second-order features contribute the most ($39.4\%$), followed by distributional features ($34.2\%$) and first-order differences ($23.7\%$). The prominence of second-order features suggests that abrupt or unnatural shifts in predictability are strong indicators of machine-generated text. While traditional distributional statistics remain informative, they are insufficient on their own. These findings support DivEye’s central claim about second-order features: modeling the evolution of surprisal yields stronger detection capabilities than relying solely on static measures. Additionally, a detailed analysis of each feature’s importance is provided in Appendix D.6.
@@ 5 Conclusion
We introduce DivEye, a lightweight classifier for AI-text detection that leverages zero-shot diversity features from token-level surprisal. Our method is model-agnostic, computationally efficient, and demonstrates strong generalization across detectors and datasets. Appendix J discusses limitations, broad impacts, and ethical considerations.
@@ Appendix A More Details on Related Work
Remark 1: Proof Sketch Consider a text sequence $X=(x_{1},x_{2},\ldots,x_{n})$ generated either by a human or by a language model $M$. The language model defines a probability distribution $P_{M}(X)=\prod_{t=1}^{n}P_{M}(x_{t}\mid x_{<t})$ where each token is chosen to maximize overall likelihood. Humans, however, produce language through a complex, multi-layered cognitive process that balances informativeness, creativity, and contextual appropriateness, rather than strictly maximizing statistical likelihood. Formally, the surprisal of token $x_{t}$ under model $M$ is defined as: $S_{M}(x_{t})=-\log P_{M}(x_{t}\mid x_{<t})$ Since $M$ is trained to assign high probability to plausible continuations, its outputs tend to minimize surprisal on average, implying that maximum likelihood generation compresses diversity: $\mathbb{E}_{X\sim P_{M}}[S_{M}(x_{t})]\leq\mathbb{E}_{X\sim P_{H}}[S_{M}(x_{t})]$ where $P_{H}$ denotes the distribution of human-generated text. Similarly, human language exhibits higher variance in surprisal due to spontaneous creative choices, idiomatic expressions, and stylistic variation, causing: $\mathrm{Var}_{X\sim P_{M}}[S_{M}(x_{t})]<\mathrm{Var}_{X\sim P_{H}}[S_{M}(x_{t})]$ We validate this theoretical intuition through empirical experiments detailed below, which confirm statistically significant differences in surprisal and diversity metrics between human-written and AI-generated texts. We collect 200 human-written essays and 200 GPT-4-Turbo-generated essays on comparable topics, provided by BiScope (Guo et al., 2024). For each essay, we computed the token-level surprisal scores using a fixed language model evaluator (GPT-2) and then calculated the mean and variance of these surprisal values per essay. Figure 8(a) shows the histogram of mean surprisal scores across the two sets, while Figure 8(b) displays the histogram of surprisal variances. The human-written texts exhibit a noticeably wider spread and heavier tails in both metrics, indicating greater unpredictability and stylistic variability. In contrast, the AI-generated essays cluster around lower mean surprisal and exhibit significantly lower variance. These results empirically confirm our theoretical claim: human language inherently reflects higher diversity and surprise, whereas AI-generated language, optimized for likelihood, tends toward more predictable and homogeneous patterns.
(a) Mean Surprisal Distribution
(b) Surprisal Variance Distribution
@@ D.7 Performance against same model
We further evaluated generalization by performing an out-of-distribution test using generations from 500 prompts drawn from the OASST (Köpf et al., 2023) and Self-Instruct (Wang et al., 2023) datasets. Table 8 reports the resulting AI detection accuracies.

### v3

@@ 1 Introduction
Contributions. We introduce DivEye  [^1] The code of our method and experiments is available at https://github.com/IBM/diveye., a lightweight classifier trained on features extracted from off-the-shelf LLMs in a zero-shot manner. These features capture diversity-based statistics of token-level surprisal, which we leverage to improve AI-text detection. Our approach focuses on capturing the distributional irregularities in AI-generated text that arise from differences in the generative process compared to human writing.
@@ 3.1 Design Hypothesis
Our hypothesis is that human-written text inherently exhibits greater stylistic diversity and unpredictability than AI-generated text. In everyday writing, humans make creative, spontaneous choices, sometimes using unexpected words or phrases, that introduce bursts of surprise amid more routine language. Our approach centers on the premise that AI-generated text, despite its fluency, often lacks the inherent diversity observed in human-written language. This divergence stems from the fundamental objective of LLMs: to maximize the likelihood of generated sequences within their learned probability distributions (Park and Choi, 2024). Consequently, AI-generated text tends to exhibit a higher degree of predictability, resulting in lower variability and surprisal compared to human-authored content. We support this hypothesis through both intuitive reasoning and empirical evidence, as detailed in  Remark 1.
Remark 1
Consider a text sequence $X=(x_{1},x_{2},\ldots,x_{n})$ generated either by a human or by a language model $M$. The language model defines a probability distribution $P_{M}(X)=\prod_{t=1}^{n}P_{M}(x_{t}\mid x_{<t})$ where each token is chosen to maximize overall likelihood.
Humans, however, produce language through a complex, multi-layered cognitive process that balances informativeness, creativity, and contextual appropriateness, rather than strictly maximizing statistical likelihood.
Formally, the surprisal of token $x_{t}$ under model $M$ is defined as:
$S_{M}(x_{t})=-\log P_{M}(x_{t}\mid x_{<t})$
Since $M$ is trained to assign high probability to plausible continuations, its outputs tend to minimize surprisal on average, implying that maximum likelihood generation compresses diversity:
$\mathbb{E}_{X\sim P_{M}}[S_{M}(x_{t})]\leq\mathbb{E}_{X\sim P_{H}}[S_{M}(x_{t})]$
where $P_{H}$ denotes the distribution of human-generated text.
Similarly, human language exhibits higher variance in surprisal due to spontaneous creative choices, idiomatic expressions, and stylistic variation, causing:
$\mathrm{Var}_{X\sim P_{M}}[S_{M}(x_{t})]<\mathrm{Var}_{X\sim P_{H}}[S_{M}(x_{t})]$
We validate this theoretical intuition through empirical experiments detailed below, which confirm statistically significant differences in surprisal and diversity metrics between human-written and AI-generated texts.
We collect 200 human-written essays and 200 GPT-4-Turbo-generated essays on comparable topics, provided by BiScope (Guo et al., 2024). For each essay, we computed the token-level surprisal scores using a fixed language model evaluator (GPT-2) and then calculated the mean and variance of these surprisal values per essay. Figure 2(a) shows the histogram of mean surprisal scores across the two sets, while Figure 2(b) displays the histogram of surprisal variances. The human-written texts exhibit a noticeably wider spread and heavier tails in both metrics, indicating greater unpredictability and stylistic variability. In contrast, the AI-generated essays cluster around lower mean surprisal and exhibit significantly lower variance. These results empirically confirm our theoretical claim: human language inherently reflects higher diversity and surprise, whereas AI-generated language, optimized for likelihood, tends toward more predictable and homogeneous patterns.
(a) Mean Surprisal Distribution
(b) Surprisal Variance Distribution
@@
These features are theoretically grounded in the Uniform Information Density (UID) hypothesis, which posits that optimal language speakers strives to distribute information uniformly Jaeger and Levy (2006). However, while human speakers aim for this efficiency, empirical evidence shows that natural language remains highly nonstationary (Piantadosi et al., 2011) and exhibits characteristic "burstiness" due to semantic shifts (Altmann et al., 2009). In contrast, LLMs, driven by likelihood-maximization objectives, often approximate the theoretical UID optimum too strictly, resulting in unnaturally flat probability curves devoid of these organic fluctuations (Holtzman et al., 2020). Consequently, we introduce first-order and second-order derivatives of surprisal, as essential metrics to quantify this ’rhythmic unpredictability’, capturing the distinct structural variance that separates human writings from LLMs.
Figure 3: Distributions of predicted class probabilities for diverse AI-text detectors. Trained and evaluated on Testbed 4 of the MAGE benchmark, DivEye shows stronger separation between human-written and Label 1 AI-generated texts, indicating greater confidence and discriminative power.
@@ 3.2 Mathematical underpinning of DivEye
Given a text sequence $X=\{x_{1},x_{2},\dots,x_{n}\}$, surprisal measures how "unexpected" each token is in context. It can be computed directly from a model’s log-probabilities, providing a principled way to quantify the local unpredictability of text.
@@
$\mathcal{D}=\{\underbrace{\mu_{s},\sigma_{s}^{2},\gamma_{1},\gamma_{2}}_{\text{Distribution}}\oplus\underbrace{\Delta\mu,\Delta\sigma^{2}}_{\text{1st-Order}}\oplus\underbrace{\sigma^{2}_{\Delta^{2}},H_{\Delta^{2}},\rho_{\Delta^{2}}}_{\text{2nd-Order}}$ (6)
@@
DivEye as a booster. Existing detectors, whether fine-tuned classifiers or zero-shot LLM-based methods, primarily rely on semantic or surface-level cues, and often falter against high-quality adversarial examples that closely mimic human writing. DivEye offers a complementary signal by capturing statistical and temporal patterns of token-level unpredictability that are orthogonal to traditional features. We integrate DivEye into both settings by augmenting detector outputs with its feature vector and training a lightweight meta-classifier (e.g., XGBoost (Chen and Guestrin, 2016) or Random Forest (Breiman, 2001)) over the combined representation. Empirically, we find that this fusion significantly boosts performance, particularly on adversarial and out-of-distribution examples, without requiring retraining or modification of the original model.
@@ 4.1 Experimental Settings
Implementation Details & Metrics. Unless stated otherwise, we use GPT-2 to compute all DivEye feature vectors. Regardless, we studied the effect of DivEye with different LLMs as base models and summarized the results in Section 4.6. In score-only detection scenarios, predictions are based solely over concatenated DivEye features. For both standalone and boosted setups, we train a lightweight XGBoost (Chen and Guestrin, 2016) classifier as a meta-model, using only DivEye features in the former, and concatenating them with the original detector’s prediction scores in the latter. Each testbed in MAGE & RAID provides predefined training and test sets, which we use for model training and evaluation. We evaluate all models using Average Accuracy (AvgAcc), AUROC, TPR@FPR=5% and F1 score to capture overall, threshold-independent, and balanced performance, respectively. For all methods, we independently optimize any required decision thresholds for AUROC and for TPR@FPR=5%, ensuring that each model is evaluated under its most favorable threshold for each metric.
@@ 4.2 Performance of DivEye
We evaluate DivEye across a wide range of challenging testbeds to assess its robustness and adaptability to both domain and model shifts. Table 1 presents the performance of DivEye on six distinct testbeds from the MAGE benchmark (Li et al., 2024): three in-distribution and three out-of-distribution. Across all testbeds, DivEye consistently achieves superior AUROC of $0.92$ on average and AvgAcc compared to existing zero-shot baselines, showcasing its ability to generalize effectively to both seen and unseen generation settings. We demonstrate that human-written and machine-generated text can be distinguished based on the hypothesis outlined in Section 3. We further include the corresponding TPR@FPR=5% performance in Table 5.
Table 4.2 benchmarks DivEye on the RAID dataset (Dugan et al., 2024), which includes a suite of diverse models, domains, attacks, and decoding strategies. DivEye outperforms a diverse set of zero-shot methods by $26.11\%$ and matches the performance of generative detection baselines, reaffirming its robustness to evasive generation strategies. Figures 4, 7 & 8 demonstrate the performance of DivEye across different domains and generator models, achieving competitive AUROC of $0.98$ and $0.93$, respectively. These results highlight DivEye’s stability and high performance across heterogeneous scenarios, underscoring its domain and model-agnostic nature.
@@
Table 2: Performance of zero-shot and open-source finetuned methods on RAID. Results are aggregated over 8 domains, 12 models, and 4 decoding strategies. $\delta$ denotes difference in AUROC from benchmark leader.
Frameworks Type TPR@ FPR=5% AUROC $\delta$
e5-small-lora [29] Fine-tuned 93.9% 0.986 -0%
DivEye (Ours) Classifier (zero-shot) 93.63% 0.984 -0.20%
Desklib AI [27] Fine-tuned 94.9% 0.973 -1.32%
SuperAnnotate [93] Fine-tuned 70.3% 0.910 -7.71%
Binoculars [45] Zero-shot 79.0% 0.844 -14.40%
RADAR [48] Fine-tuned 65.6% 0.819 -16.94%
GLTR [38] Zero-shot 59.7% 0.727 -26.27%
@@ 4.3 Robustness to Adversarial Attacks and Multilingual Text
Table 3: Performance of DivEye and baselines on adversarial benchmarks, MAGE & RAID. - RAID reports TPR@FPR=5% instead of AvgAcc.
Settings Methods AvgAcc AUROC [MAGE] Testbed 8: Paraphrasing Attack Paraphrased via GPT-3.5-Turbo Longformer [11] 69.34% 0.76 BiScope [44] 69.30% 0.81 DivEye (Ours) 76.49% 0.87 [RAID] Adversarial Attacks Paraphrase, Whitespace, Misspelling, Homoglyph, Article Deletion & more Desklib AI [27] 91.2% 0.948 e5-small-lora [29] 85.7% 0.968 DivEye (Ours) 80.52% 0.951 Binoculars [45] 69.32% - RADAR [48] 63.9% 0.819 GLTR [38] 51.5% 0.709
@@ 4.5 Effectiveness of Boosting by DivEye
We empirically verify that DivEye-based diversity features can act as performance boosters for a wide range of detection models. To integrate DivEye, we concatenate its feature vector with the original model’s prediction scores and train a lightweight XGBoost classifier as a meta-model.
Table 4 illustrates improvements in AUROC and AvgAcc when diversity metrics are appended to existing frameworks such as RADAR, Binoculars, DetectLLM, BiScope and FastDetectGPT. Across all evaluated baselines, the inclusion of diversity features consistently leads to better detection scores by over $18.7\%$. Additionally, existing frameworks in combination with DivEye demonstrate substantial performance gains when evaluated against paraphrasing attacks. This validates the hypothesis that static and dynamic surprisal-based features capture orthogonal information to traditional heuristics, making them a valuable addition to any detection pipeline. We also investigate how much DivEye contributes to the final prediction relative to the underlying detector. Appendix D.4 reports a detailed feature-importance analysis, highlighting the complementary signals captured by DivEye within the boosted ensembles.
@@ 4.6 Ablation Studies
DivEye’s Performance on Different Base Models. To assess the adaptability of DivEye across different language model backbones, we evaluate its performance, on Testbed 4 of the MAGE benchmark, when instantiated with various base LLMs used to compute token-level surprisal. As shown in Figure 6(a), DivEye consistently performs well across all models, achieving an AUROC of $0.88$ with GPT2, $0.91$ with Llama-3.1-8B, and $0.90$ with Falcon-7B. Notably, even the smallest model, GPT2, performs competitively, and human classification accuracy improves with larger models, suggesting that higher-capacity LMs better capture stylistic diversity. These results highlight DivEye’s robustness and efficiency across scales, making it suitable for resource-constrained settings. Appendix D.3 further reports baseline and DivEye performance across various base models.
Relevance of DivEye’s Features. DivEye’s feature set (Equation 6) captures token-level surprisal patterns across multiple orders, including distributional moments, first-order shifts, and second-order dynamics. To assess the contribution of each group, we compute feature importances from a trained XGBoost model. On average, second-order features contribute the most ($39.4\%$), followed by distributional features ($34.2\%$) and first-order differences ($23.7\%$). The prominence of second-order features suggests that abrupt or unnatural shifts in predictability are strong indicators of machine-generated text. While traditional distributional statistics remain informative, they are insufficient on their own. These findings support DivEye’s central claim about second-order features: modeling the evolution of surprisal yields stronger detection capabilities than relying solely on static measures. Additionally, Appendix D.6 contains a detailed discussion about the contribution of each component in DivEye, including both feature-importance estimates and statistical leave-one-out ablation results.
@@ 5 Conclusion
We introduce DivEye, a lightweight classifier for AI-text detection that leverages zero-shot diversity features from token-level surprisal. Our method is model-agnostic, computationally efficient, and demonstrates strong generalization across detectors and datasets. Appendix J provides a detailed discussion of the method’s limitations, broader impacts, and associated ethical considerations.
@@ Appendix D Additional Results
Table 5: TPR@FPR=5% on Testbed 8 (Unseen Domains & Models).
LogRank Entropy DetectLLM FastDetect Binoculars BiScope DivEye
TPR 47.1 30.1 51.5 76.3 68.7 99.8 99.6
@@ D.7 Performance against same model
We further evaluated generalization by performing an out-of-distribution test using generations from 500 prompts drawn from the OASST (köpf2023openassistantconversationsdemocratizing) and Self-Instruct (Wang et al., 2023) datasets. Table 9 reports the resulting AI detection accuracies.
@@ D.8 Performance against Longformer
D.9 Performance across edge-cases
D.9.1 Performance across non-native-written texts
To evaluate potential bias against non-native speakers, a known vulnerability in many perplexity-based detectors, we conducted a targeted evaluation using the COREFL dataset (Lozano et al., 2020). This dataset consists of 1426 essays written by native German and Spanish speakers, explicitly categorized by their English proficiency levels (ranging from A1 to C2).
We trained (or configured) DivEye, Binoculars and FastDetectGPT using a balanced set comprising 75% of the COREFL human texts and an equivalent number of AI-generated samples drawn from MAGE Testbed 4. We evaluate performance on the held-out 25% to measure detection accuracy across proficiency brackets.
Table 11: Detection Accuracy on Non-Native English Writing.
Framework Average Accuracy
Binoculars 68.34%
FastDetectGPT 71.71%
DivEye (Ours) 82.07%
As shown in Table 11, DivEye significantly outperforms existing baselines in identifying non-native human text, achieving a robust 82.1% average accuracy compared to 68.34% for Binoculars and 71.7% for FastDetectGPT.
Figure 12: Granular depiction of various frameworks performance on various non-native texts.
Crucially, as illustrated in Figure 12, our granular analysis reveals that performance of all frameworks correlates with linguistic complexity: A1 texts (the most beginner class) are the most difficult to detect, while accuracy steadily improves as proficiency rises. As for DivEye, this suggests that detection becomes easier as the text exhibits the richer rhythmic signatures of advanced writing.
D.9.2 Performance across legal & professional documents
A common criticism in AI text detectors is their potential fragility when applied to highly formal human writing, such as legal contracts or technical specifications. In these domains, human writers often strive for low entropy and rigid structure, traits typically associated with machine generation. To address this concern, we evaluated DivEye on the LEDGAR dataset (Tuggener et al., 2020), a corpus of legal provisions extracted from US SEC filings.
Table 12: Detection Accuracy on LEDGAR dataset.
Framework Accuracy
Binoculars 91.67%
FastDetectGPT 94.8%
DivEye (Ours) 99.49%
We adopted a robust evaluation protocol using a 75-25 train-test split to ensure generalization. As shown in Table 12, despite the constrained and repetitive nature of contract language, DivEye successfully outperforms other frameworks on this task. These findings provide empirical evidence that DivEye’s core hypothesis holds even in challenging edge cases.
D.9.3 Visualizing texts under observation
Analyzing the probability distributions across domains reveals a distinct hierarchy of detectability. We have already seen a distribution in Figure 3 for DivEye’s effectiveness on MAGE Testbed 4.
Contrary to the concern that formal writing might resemble algorithmic generation, we find that in Figure 13 professional legal text (LEDGAR) is the most easily distinguishable from AI, with its probability distribution shifted furthest towards the human extreme.
Figure 13: Predicted distributions of predicted class probabilities for diverse human written texts. We utilize the LEDGAR, COREFL and the MAGE Testbed 8 texts to plot this distribution.
This suggests that while professional writing is structured, it contains sharp, domain-specific peaks in information density that are representative of human writing. Furthermore, Standard (MAGE Testbed 8) and Non-Native (COREFL) texts exhibit remarkably similar distributions, effectively overlapping in the high-probability region. We hope to investigate this further in our future works.
D.9.4 Text sensitivity analysis of DivEye
To determine the minimum linguistic context required for reliable detection, we evaluated the performance of DivEye across varying text lengths. We truncated test samples to specific token counts ranging from $N=10$ to $N=256$ and measured the AUROC at each interval. As illustrated in Figure 14, performance follows a predictable scaling law relative to information availability.
Detection is constrained on extremely short fragments, yielding an AUROC of $0.62$ at 10 tokens; this is expected, however, the system demonstrates rapid convergence: performance improves steadily as the context window expands, reaching a robust AUROC of $0.88$ by 256 tokens. This confirms that while DivEye benefits from longer contexts, it effectively captures the structural signature of human writing within standard paragraph lengths.
We discuss short length texts as a limitation in Appendix J.
Figure 14: DivEye’s performance on different text lengths (characterized by tokens) on MAGE Testbed 4.